A string is stretched to a tension of 100 N, and has a linear density of...

  1. A string is stretched to a tension of 100 N, and has a linear density of 0.025 kg/m. An input disturbance causing a sinusoidal wave has a frequency of 150 Hz, with an amplitude of 5 cm.

    1. Determine the speed of the wave.

    2. Determine the wavelength.

    3. Write down the equation describing the displacement of the string as a function of

      the position and time.
Answer #1

a)

v^2 = T/u = 100 / 0.025

v = 63.246 m/s

B)

L = v / f = 63.246 / 150 = 0.4216 m

C)

x = A cos ( wt)

x = 5 cos ( 2 pi * 150 t)

x = 5 cos ( 942 t)

======

Comment before rate in case any doubt, will reply for sure.. Goodluck